OpenStudy (some.random.cool.kid):

Solve the equation 4x2 + 8x + 1 = 0 by completing the square.

OpenStudy (some.random.cool.kid):

@Hero @dan815 @iambatman

OpenStudy (some.random.cool.kid):

@phi @amistre64

OpenStudy (some.random.cool.kid):

@Nnesha mind helping :D

OpenStudy (some.random.cool.kid):

not to much this time...

OpenStudy (some.random.cool.kid):

@Nnesha

OpenStudy (anonymous):

First, make the coefficient of the x^2 term 1 by dividing both sides by 4.

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

so 4 by which sides?

OpenStudy (anonymous):

Divide all terms on both sides of the equation by 4.

OpenStudy (some.random.cool.kid):

the x^2 and the 8x by 4?

OpenStudy (anonymous):

Everything, yes. What will it look like?

OpenStudy (some.random.cool.kid):

uh... well 8x would be 2x?

OpenStudy (anonymous):

Yup. Keep going.

OpenStudy (some.random.cool.kid):

and what about the ^2

OpenStudy (some.random.cool.kid):

is that 4^2?

OpenStudy (anonymous):

Divide 4x^2 by 4. What do you get?\[\frac{ 4x^2 }{ 4 } = ?\]

OpenStudy (some.random.cool.kid):

just x^2?

OpenStudy (anonymous):

Right. Keep going.

OpenStudy (some.random.cool.kid):

1 divided by 4 is that 4?

OpenStudy (anonymous):

No. It's \(\frac{1}{4}\). What about the right hand side? Gotta do that too.

OpenStudy (some.random.cool.kid):

0

OpenStudy (anonymous):

Good. So, dividing everything by 4 gives\[x^2 + 2x + \frac{ 1 }{ 4 } = 0\]OK so far?

OpenStudy (some.random.cool.kid):

alright

OpenStudy (anonymous):

Now, we don't want the 1/4 on the left hand side, so subtract 1/4 from both sides. What will it look like then?

OpenStudy (some.random.cool.kid):

uh - 1/4?

OpenStudy (some.random.cool.kid):

thats a negative

OpenStudy (anonymous):

OK. What does the whole equation look like now?

OpenStudy (some.random.cool.kid):

x^2 + 2x - 1/4?

OpenStudy (anonymous):

Not quite. What happened to the equals sign?

OpenStudy (some.random.cool.kid):

idk is it x^2 + 2x - 1/4 =0 or x^2 + 2x = 1/4?

OpenStudy (anonymous):

We better back up a step. You had\[x^2 + 2x + \frac{ 1 }{ 4 } = 0\]To get rid of the 1/4 on the left hand side, you need to subtract it. And the rules of equality say that whatever you do one side of the equation you have to do to the other side as well. So you have to subtract 1/4 from both sides. So\[x^2 + 2x + \frac{ 1 }{ 4 } - \frac{ 1 }{ 4 } = 0-\frac{ 1 }{ 4 }\]Simplify both sides. What will it look like?

OpenStudy (some.random.cool.kid):

wait but if you canceled the other 1/4ths and your left - 1/4 then what is left i dont get it all i se is x^2 + 2x - 1/4

OpenStudy (anonymous):

This is not an equation. What happened to the equal sign. It can't just disappear.

OpenStudy (some.random.cool.kid):

you dont have one you used it when you crossed out the 1/4ths

OpenStudy (some.random.cool.kid):

thats how i was tought...

OpenStudy (anonymous):

Sorry, no. Equals signs don't disappear. What you are left with is\[x^2 + 2x = -\frac{ 1 }{ 4 }\]Do you understand?

OpenStudy (some.random.cool.kid):

ooo ok

OpenStudy (anonymous):

OK. So now you're ready to complete the square on the left hand side. Do you remember how to do it?

OpenStudy (some.random.cool.kid):

some more dividing? lawl

OpenStudy (anonymous):

Not really :) Do you know how to complete the square?

OpenStudy (some.random.cool.kid):

mm take half the sq on both sides?

OpenStudy (anonymous):

Something like that. You need to add a number to the left hand side that will make it a perfect square trinomial. Look at the coefficient of the 'x' term. Take half of that coefficient and then square it. What number do you get?

OpenStudy (some.random.cool.kid):

your left with x +2? after taking away one x from each of the sides?

OpenStudy (anonymous):

No. What is the coefficient of the 'x' term on the left hand side?

OpenStudy (some.random.cool.kid):

sorry idk that..

OpenStudy (anonymous):

What number is multiplying x on the left hand side?

OpenStudy (some.random.cool.kid):

1

OpenStudy (anonymous):

Oh my. The equation you're working with now is\[x^2 + 2x = -\frac{ 1 }{ 4 }\]The 'x' term on the left hand side is \(2x\). The coefficient is the number that is multiplying the \(x\). What number is it?

OpenStudy (some.random.cool.kid):

omg i said idk geez...

OpenStudy (anonymous):